OpenStudy (anonymous):

Solve for x: x-4 ------- = 12 3 A) 12 B) 28 C) 23 D) 20 44

OpenStudy (anonymous):

none of the above.

OpenStudy (anonymous):

(x-4) / 3 = 12 multiply both sides by 3: (x-4) = 36 hence, x - 4 = 36 solve for x and get your answer ps.: as @dhatraditya said, NOA

OpenStudy (anonymous):

im confused lol

OpenStudy (anonymous):

where?

OpenStudy (anonymous):

okay hold on, which grade are you in?

OpenStudy (anonymous):

im talking about the anwer being none of the above

OpenStudy (anonymous):

its 44

OpenStudy (anonymous):

then you must've typed something wrong when posting the question